# Gravity from (A)dS Yang-Mills: a natural reduction

### v2

@@ 2.1 Differential identities implied by $\mathcal{D}_{\Omega}\mathcal{F}=0$
$\mathcal{D}_{\Omega}\mathcal{F}=d\mathcal{F}+\left[\Omega,\mathcal{F}\right]=0$ (2.13)
@@
$\begin{split}\mathcal{D}_{\Omega}\mathcal{F}&=\frac{1}{2}\left(d\mathfrak{R}^{ab}+\frac{1}{2}f^{ab}_{ij,cd}\varpi^{ij}\wedge\mathfrak{R}^{cd}+\lambda\left(d\mathfrak{E}^{ab}+\frac{1}{2}f^{ab}_{cd,ef}\varpi^{cd}\wedge\mathfrak{E}^{ef}+2K^{ab}_{c,d}\vartheta^{c}\wedge\mathfrak{T}^{d}\right)\right)J_{ab}\\ &+\sqrt{\frac{\lambda}{{\alpha}}}\left(d\mathfrak{T}^{a}+\frac{1}{2}f^{a}_{cd,b}\varpi^{cd}\wedge\mathfrak{T}^{b}+f^{a}_{b,cd}\vartheta^{b}\wedge\frac{1}{2}\mathfrak{R}^{cd}+\frac{\lambda}{2}\left(f^{a}_{b,cd}\vartheta^{b}\wedge\mathfrak{E}^{cd}\right)\right)P_{a}\end{split}$
@@
$\frac{1}{2}\left(\mathcal{D}_{\varpi}\mathfrak{R}^{ab}\right)J_{ab}+\sqrt{\frac{\lambda}{{\alpha}}}\left(\mathcal{D}_{\varpi}\mathfrak{T}^{a}-\vartheta^{b}\wedge\mathfrak{R}^{a}\,_{b}\right)P_{a}=0\,$ (2.15)
that is
$\mathcal{D}_{\varpi}\mathfrak{R}^{ab}$ $=$ 00 0 (2.16)
$\mathcal{D}_{\varpi}\mathfrak{T}^{a}-\vartheta^{b}\wedge\mathfrak{R}^{a}\,_{b}$ $=$ 00 0 (2.17)
@@ 2.2 Transformation properties of connection one-form
In order to establish the desired interpretation of the gauge connection, it is essential that its components transform correctly under the action of the local Lorentz group in the limit where the full group reduces to the Poincaré group. To check this, let us first consider gauge transformations of the gauge connection $\Omega$. We have
@@
$\begin{split}\Omega^{\prime}=\Omega+\varepsilon^{ab}[\Omega,J_{ab}]+\varepsilon^{a}[\Omega,P_{a}]+d\varepsilon^{ab}J_{ab}+d\varepsilon^{a}P_{a}\end{split}$ (2.19)
@@
$\begin{split}\frac{1}{2}\delta\varpi^{ef}&=d\varepsilon^{ef}+f_{ab,cd}^{ef}\frac{1}{2}\varpi^{ab}\varepsilon^{cd}\\ \delta\vartheta^{e}&=f_{a,cd}^{e}\vartheta^{a}\varepsilon^{cd}\end{split}$ (2.22)
Hence, the component $\varpi$ of the connection transforms non-homogeneously under Lorentz transformations, just like the Lorentz connection, while the component $\vartheta$ transforms like a Lorentz vector, similarly to tetrad one-forms. In other words, in the limit $\alpha\rightarrow 0$, $\vartheta$ is no longer a gauge connection, whereas $\varpi$ is, and the residual gauge group is the Lorentz group.
@@ 3 Euler-Lagrange equations of YM theory
$\frac{\delta}{\delta\varpi_{ab}}S_{YM}$ $=$ $\frac{\hbar}{4g^{2}}\left[\mathcal{D}_{\varpi}(*\mathfrak{R}^{ab})+\lambda\mathcal{D}_{\varpi}(*\mathfrak{E}^{ab})+\lambda(\vartheta^{a}\wedge*\mathfrak{T}^{b}-\vartheta^{b}\wedge*\mathfrak{T}^{a})\right]$ (3.3)
We observe that the dynamics of the curvature $\mathfrak{R}$ in (3.2) is characterised by a term related to the covariant derivative of the two-form $\mathfrak{T}$, while the equation (3.3) for the two-form $\mathfrak{T}$ is characterised by the presence of the derivatives of the curvature $\mathfrak{R}$, as well as by a linear term in $\mathfrak{T}$. These terms arise directly from the presence of the Euler classes in the action (2.10).
In order to include sources in the Yang-Mills equations, given a gauge-invariant source action, we define the 3-forms :
$\Theta^{a}=\frac{\delta S_{sources}}{\delta\vartheta_{a}}$ (3.4)
@@
$\delta S_{source}=\int_{\mathcal{M}}\left(\delta\Omega\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}J\right)\,$ (3.9)
@@
$\frac{\hbar}{2g^{2}}\,\mathcal{D}_{\Omega}*\mathcal{F}$ $=$ $-J$ (3.11)
@@
$\mathcal{D}_{\varpi}(*\mathfrak{T}^{a})+\vartheta_{b}\wedge*\mathfrak{R}^{ab}+\lambda\vartheta_{b}\wedge*\mathfrak{E}^{ab}$ $=$ $\chi\Theta^{a}$ (3.12)
$\mathcal{D}_{\varpi}(*\mathfrak{R}^{ab})+\lambda\mathcal{D}_{\varpi}(*\mathfrak{E}^{ab})+\lambda(\vartheta^{a}\wedge*\mathfrak{T}^{b}-\vartheta^{b}\wedge*\mathfrak{T}^{a})$ $=$ $\lambda\chi{\Sigma}^{ab}$ (3.13)
@@
$+\left(\frac{1}{\sqrt{\alpha\lambda}}\left[d\Theta^{a}+\frac{1}{2}f^{a}_{cd,b}\varpi^{cd}\wedge\Theta^{b}\right]+\frac{1}{2}\sqrt{\frac{\lambda}{\alpha}}f^{a}_{c,bd}\vartheta^{c}\wedge\Sigma^{bd}\right)P_{a}$
@@
$\mathcal{D_{\varpi}}\Sigma^{ab}+\vartheta^{a}\wedge\Theta^{b}-\vartheta^{b}\wedge\Theta^{a}$ $=$ 00 0 (3.15)
$\mathcal{D}_{\varpi}\Theta^{a}+\lambda\vartheta_{b}\wedge{\Sigma}^{ab}$ $=$ 00 0 (3.16)
Eqs. (3.15) and (3.16) are similar in form to the conservation laws obtained in the Einstein-Cartan-Sciama-Kibble (ECSK) approach [7, 29]. However, in our case they stem from the gauge conservation of the three-form $J$ imposed by Yang-Mills theory.
@@
$\mathcal{D}_{\varpi}*\mathcal{F}^{ab}=0$ (3.17)
that is $\mathcal{D}_{\varpi}(*\mathfrak{R}^{ab})+\lambda\mathcal{D}_{\varpi}(*\mathfrak{E}^{ab})=0$, the Eqs. (3.12), (3.13) result in
$\mathcal{D}_{\varpi}(*\mathfrak{T}^{a})+\vartheta_{b}\wedge*\mathfrak{R}^{ab}+\lambda\vartheta_{b}\wedge*\mathfrak{E}^{ab}$ $=$ $\chi\Theta^{a}$ (3.18)
$\vartheta^{a}\wedge*\mathfrak{T}^{b}-\vartheta^{b}\wedge*\mathfrak{T}^{a}$ $=$ $\chi{\Sigma}^{ab}$ (3.19)
In particular, in this case, the equations of motion describe a non-propagating torsion field $\mathfrak{T}$, due to the algebraic relationship between torsion and spin that is consequently established.
By inspecting the Eqs. (3.12) and (3.13) it is conceivable that different choices besides $\mathcal{D}_{\varpi}*\mathcal{F}^{ab}=0$ can lead to a consistent reduction of the dynamics compatible with gravity. We shall investigate these alternatives in future work.
@@ 4 Geometric character of the gauge fields for $\alpha\to 0$
So far, we have considered a standard Yang-Mills theory on Minkowski space-time with structure group ${G}_{\alpha}$ and a Yang-Mills action given by (2.8), that is invariant under global Poincaré transformations. The connection one-form $\Omega$ decomposes along the Lie algebra generators in terms of $\vartheta$ and $\varpi$, which transform appropriately under the action of the gauge group. Moreover they behave as covariant vectors under global Poincaré transformations.
In Section 2.2, we remarked that the $P$-gauge translations no longer encode the gauge transformations of $\vartheta$ in the limit $\alpha\to 0$ (see Eqn. (2.22)). We are going to show that the inhomogeneous part of the $P$-gauge transformations can be reinterpreted in terms of diffeomorphisms. Such identification will lead to the interpretation of the gauge field $\vartheta$ as the tetrad one form in the limit $\alpha\rightarrow 0$.
@@ 4.1 From gauge translations to diffeomorphisms: soldering
::=:= $\delta_{P(\varepsilon^{c})}(\vartheta_{\mu}^{a})+\sqrt{\frac{\lambda}{\alpha}}\delta_{J(\varepsilon^{cd})}(\vartheta_{\mu}^{a})$
@@
::=:= $\delta_{J(\varepsilon^{cd})}(\frac{1}{2}\varpi_{\mu}^{ab})+\sqrt{\alpha\lambda}K^{ab}_{d,c}\vartheta_{\mu}^{d}\varepsilon^{c}$
where we denoted by $\varepsilon^{a}$ and $\varepsilon^{ab}$ the coordinate-dependent parameters of the gauge transformations [^5] Hereafter, $\delta$ without subscripts (GCT) or (CGCT) will always refer to gauge transformations, and their parameters, functions of spacetime coordinates, will be enclosed in parentheses. and we have singled out the different contributions along the generators of the Lie algebra. At the same time, the connection transforms infinitesimally under general coordinate transformation (GCT) $y^{\mu}(x)=x^{\mu}+\xi^{\mu}(x)$, according to
$\delta_{GCT(\xi)}\Omega^{I}_{\mu}=\mathcal{L}_{\xi}\Omega^{I}_{\mu}=\xi^{\nu}\partial_{\nu}\Omega_{\mu}^{I}+\Omega_{\nu}^{I}\partial_{\mu}\xi^{\nu}$ (4.3)
@@
$\begin{split}\delta_{GCT(\xi)}\vartheta^{a}_{\mu}&=\xi^{\nu}\partial_{\nu}\vartheta_{\mu}^{a}+\vartheta_{\nu}^{a}\partial_{\mu}\xi^{\nu}\\ \delta_{GCT(\xi)}\varpi^{ab}_{\mu}&=\xi^{\nu}\partial_{\nu}\varpi_{\mu}^{ab}+\varpi_{\nu}^{ab}\partial_{\mu}\xi^{\nu}\end{split}$ (4.4)
In order to study the interplay of gauge and coordinate transformations, it is convenient to consider a covariant generalization of the latter (CGCT) [30]. For the full gauge potential this is defined as
@@
$\begin{split}\delta_{CGCT(\xi)}(\frac{1}{2}\varpi^{ab}_{\mu})&=\delta_{GCT(\xi)}(\frac{1}{2}\varpi^{ab}_{\mu})\\ &-\left(\partial_{\mu}(\xi^{\rho}\frac{1}{2}\varpi^{ab}_{\rho})+(\xi^{\rho}\frac{1}{2}\varpi^{cd}_{\rho})\frac{1}{2}\varpi^{ef}_{\mu}f^{ab}_{cd,ef}+(\xi^{\rho}\vartheta^{c}_{\rho})\vartheta^{d}_{\mu}\lambda K^{ab}_{c,d}\right)\end{split}$ (4.7)
@@
$\begin{split}\delta_{CGCT(\xi)}\vartheta^{a}_{\mu}&=\delta_{GCT(\xi)}\vartheta^{a}_{\mu}-\delta_{P(\xi^{\rho}\vartheta^{c}_{\rho})}(\vartheta_{\mu}^{a})-\delta_{J(\xi^{\rho}\varpi^{cd}_{\rho})}(\vartheta_{\mu}^{a})\\ \delta_{CGCT(\xi)}\varpi^{ab}_{\mu}&=\delta_{GCT(\xi)}\varpi^{ab}_{\mu}-\delta_{J(\xi^{\rho}\frac{1}{2}\varpi^{cd}_{\rho})}(\varpi_{\mu}^{ab})-2\lambda\xi^{\rho}\vartheta^{c}_{\rho}\vartheta_{\mu}^{d}K_{c,d}^{ab}\end{split}$ (4.8)
@@
$\begin{split}\delta_{CGCT(\xi)}\vartheta^{a}_{\mu}=\xi^{\rho}\mathfrak{T}^{a}_{\rho\mu}\end{split}$ (4.9)
@@
$\begin{split}\delta_{CGCT(\xi)}\varpi^{ab}_{\mu}=\xi^{\rho}\mathfrak{R}^{ab}_{\rho\mu}-2\lambda\xi^{\rho}\vartheta^{c}_{\rho}\vartheta_{\mu}^{d}K_{c,d}^{ab}\end{split}$ (4.10)
@@
$\begin{split}\delta_{GCT(\xi)}\vartheta^{a}_{\mu}-\delta_{P(\xi^{\rho}\vartheta^{c}_{\rho})}(\vartheta_{\mu}^{a})-\delta_{J(\xi^{\rho}\varpi^{cd}_{\rho})}(\vartheta_{\mu}^{a})&=\xi^{\rho}\mathfrak{T}^{a}_{\rho\mu}\\ \delta_{GCT(\xi)}\varpi^{ab}_{\mu}-\delta_{J(\xi^{\rho}\frac{1}{2}\varpi^{cd}_{\rho})}(\varpi_{\mu}^{ab})&=\xi^{\rho}\mathfrak{R}^{ab}_{\rho\mu}\end{split}$ (4.11)
As anticipated, CGCT allow to rewrite the diffeomorphisms in terms of gauge transformations and covariant tensors. The first of Eqs. (4.11), appropriately rearranged, gives the soldering equation
@@
$\begin{cases}\vartheta^{a}\rightarrow\theta^{a}\,\,\,\text{tetrad 1-form}\\ \varpi^{ab}\rightarrow\omega^{ab}\,\,\,\text{Lorentz connection 1-form}\end{cases}$ (4.13)
Consequently, by recalling the definition of the fields in Eqs. (2.5) and (2.6), we can give a consistent geometric interpretation to the components of the curvature 2-form $\mathcal{F}$:
$\begin{cases}\mathfrak{T}^{a}\rightarrow T^{a}\,\,\,\text{ torsion 2-form }\\ \mathfrak{R}^{ab}\rightarrow R^{ab}\,\,\,\text{Lorentz curvature 2-form }\end{cases}$ (4.14)

### v3

@@ 2.1 Differential identities implied by $\mathcal{D}_{\Omega}\mathcal{F}=0$
$\mathcal{D}_{\Omega}\mathcal{F}=d\mathcal{F}+\left[\Omega,\mathcal{F}\right]=0,$ (2.13)
@@
$\begin{split}\mathcal{D}_{\Omega}\mathcal{F}&=\frac{1}{2}\left(d\mathfrak{R}^{ab}+\frac{1}{2}f^{ab}_{ij,cd}\varpi^{ij}\wedge\mathfrak{R}^{cd}+\lambda\left(d\mathfrak{E}^{ab}+\frac{1}{2}f^{ab}_{cd,ef}\varpi^{cd}\wedge\mathfrak{E}^{ef}+2K^{ab}_{c,d}\vartheta^{c}\wedge\mathfrak{T}^{d}\right)\right)J_{ab}\\ &+\sqrt{\frac{\lambda}{{\alpha}}}\left(d\mathfrak{T}^{a}+\frac{1}{2}f^{a}_{cd,b}\varpi^{cd}\wedge\mathfrak{T}^{b}+f^{a}_{b,cd}\vartheta^{b}\wedge\frac{1}{2}\mathfrak{R}^{cd}+\frac{\lambda}{2}\left(f^{a}_{b,cd}\vartheta^{b}\wedge\mathfrak{E}^{cd}\right)\right)P_{a}.\end{split}$
@@
$\frac{1}{2}\left(\mathcal{D}_{\varpi}\mathfrak{R}^{ab}\right)J_{ab}+\sqrt{\frac{\lambda}{{\alpha}}}\left(\mathcal{D}_{\varpi}\mathfrak{T}^{a}-\vartheta^{b}\wedge\mathfrak{R}^{a}\,_{b}\right)P_{a}=0,$ (2.15)
that is,
$\mathcal{D}_{\varpi}\mathfrak{R}^{ab}$ $=$ $0$ (2.16)
$\mathcal{D}_{\varpi}\mathfrak{T}^{a}-\vartheta^{b}\wedge\mathfrak{R}^{a}\,_{b}$ $=$ $0.$ (2.17)
@@ 2.2 Transformation properties of connection one-form
To establish the desired interpretation of the gauge connection, it is essential that its components transform correctly under the action of the local Lorentz group in the limit where the full group reduces to the Poincaré group. To check this, let us first consider gauge transformations of the gauge connection $\Omega$. We have
@@
$\begin{split}\Omega^{\prime}=\Omega+\varepsilon^{ab}[\Omega,J_{ab}]+\varepsilon^{a}[\Omega,P_{a}]+d\varepsilon^{ab}J_{ab}+d\varepsilon^{a}P_{a}.\end{split}$ (2.19)
@@
$\begin{split}\frac{1}{2}\delta\varpi^{ef}&=d\varepsilon^{ef}+f_{ab,cd}^{ef}\frac{1}{2}\varpi^{ab}\varepsilon^{cd}\\ \delta\vartheta^{e}&=f_{a,cd}^{e}\vartheta^{a}\varepsilon^{cd}.\end{split}$ (2.22)
Hence, the component $\varpi$ of the connection transforms nonhomogeneously under Lorentz transformations, just like the Lorentz connection, while the component $\vartheta$ transforms like a Lorentz vector, similarly to tetrad one-forms. In other words, in the limit $\alpha\rightarrow 0$, $\vartheta$ is no longer a gauge connection, whereas $\varpi$ is, and the residual gauge group is the Lorentz group.
@@ 3 Euler-Lagrange equations of YM theory
$\frac{\delta}{\delta\varpi_{ab}}S_{YM}$ $=$ $\frac{\hbar}{4g^{2}}\left[\mathcal{D}_{\varpi}(*\mathfrak{R}^{ab})+\lambda\mathcal{D}_{\varpi}(*\mathfrak{E}^{ab})+\lambda(\vartheta^{a}\wedge*\mathfrak{T}^{b}-\vartheta^{b}\wedge*\mathfrak{T}^{a})\right].$ (3.3)
We observe that the dynamics of the curvature $\mathfrak{R}$ in (3.2) is characterized by a term related to the covariant derivative of the two-form $\mathfrak{T}$, while Eq (3.3) for the two-form $\mathfrak{T}$ is characterized by the presence of the derivatives of the curvature $\mathfrak{R}$, as well as by a linear term in $\mathfrak{T}$. These terms arise directly from the presence of the Euler classes in the action (2.10).
To include sources in the Yang-Mills equations, given a gauge-invariant source action, we define the three-forms as
$\Theta^{a}=\frac{\delta S_{sources}}{\delta\vartheta_{a}},$ (3.4)
@@
$\delta S_{source}=\int_{\mathcal{M}}\left(\delta\Omega\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}J\right).$ (3.9)
@@
$\frac{\hbar}{2g^{2}}\,\mathcal{D}_{\Omega}*\mathcal{F}=-J,$ (3.11)
@@
$\mathcal{D}_{\varpi}(*\mathfrak{T}^{a})+\vartheta_{b}\wedge*\mathfrak{R}^{ab}+\lambda\vartheta_{b}\wedge*\mathfrak{E}^{ab}$ $=$ $\chi\Theta^{a}\,$, (3.12)
$\mathcal{D}_{\varpi}(*\mathfrak{R}^{ab})+\lambda\mathcal{D}_{\varpi}(*\mathfrak{E}^{ab})+\lambda(\vartheta^{a}\wedge*\mathfrak{T}^{b}-\vartheta^{b}\wedge*\mathfrak{T}^{a})$ $=$ $\lambda\chi{\Sigma}^{ab}\,$, (3.13)
@@
$+\left(\frac{1}{\sqrt{\alpha\lambda}}\left[d\Theta^{a}+\frac{1}{2}f^{a}_{cd,b}\varpi^{cd}\wedge\Theta^{b}\right]+\frac{1}{2}\sqrt{\frac{\lambda}{\alpha}}f^{a}_{c,bd}\vartheta^{c}\wedge\Sigma^{bd}\right)P_{a}.$
@@
$\mathcal{D_{\varpi}}\Sigma^{ab}+\vartheta^{a}\wedge\Theta^{b}-\vartheta^{b}\wedge\Theta^{a}$ $=$ $0,$ (3.15)
$\mathcal{D}_{\varpi}\Theta^{a}+\lambda\vartheta_{b}\wedge{\Sigma}^{ab}$ $=$ $0.$ (3.16)
Equations. (3.15) and (3.16) are similar in form to the conservation laws obtained in the Einstein-Cartan-Sciama-Kibble (ECSK) approach [Krasnov:2020lku, PhysRevD.21.3269]. However, in our case they stem from the gauge conservation of the three-form $J$ imposed by Yang-Mills theory.
@@
$\mathcal{D}_{\varpi}*\mathcal{F}^{ab}=\mathcal{D}_{\varpi}(*\mathfrak{R}^{ab})+\lambda\mathcal{D}_{\varpi}(*\mathfrak{E}^{ab})=0,$ (3.17)
Eqs. (3.12)-and (3.13) result in
$\mathcal{D}_{\varpi}(*\mathfrak{T}^{a})+\vartheta_{b}\wedge*\mathfrak{R}^{ab}+\lambda\vartheta_{b}\wedge*\mathfrak{E}^{ab}$ $=$ $\chi\Theta^{a},$ (3.18)
$\vartheta^{a}\wedge*\mathfrak{T}^{b}-\vartheta^{b}\wedge*\mathfrak{T}^{a}$ $=$ $\chi{\Sigma}^{ab}.$ (3.19)
In particular, in this case, the equations of motion describe a nonpropagating torsion field $\mathfrak{T}$, due to the algebraic relationship between torsion and spin that is consequently established.
By inspecting Eqs. (3.12) and (3.13) it is conceivable that different choices besides $\mathcal{D}_{\varpi}*\mathcal{F}^{ab}=0$ can lead to a consistent reduction of the dynamics compatible with gravity. We shall investigate these alternatives in future work.
@@ 4 Geometric character of the gauge fields for $\alpha\to 0$
So far, we have considered a standard Yang-Mills theory on Minkowski spacetime with structure group ${G}_{\alpha}$ and a Yang-Mills action given by (2.8), that is invariant under global Poincaré transformations. The connection one-form $\Omega$ decomposes along the Lie algebra generators in terms of $\vartheta$ and $\varpi$, which transform appropriately under the action of the gauge group. Moreover they behave as covariant vectors under global Poincaré transformations.
In Sec. 2.2, we remarked that the $P$-gauge translations no longer encode the gauge transformations of $\vartheta$ in the limit $\alpha\to 0$ [see Eqn. (2.22)]. We are going to show that the inhomogeneous part of the $P$-gauge transformations can be reinterpreted in terms of diffeomorphisms.
@@ 4.1 From gauge translations to diffeomorphisms: Soldering
$:=$ $\delta_{P(\varepsilon^{c})}(\vartheta_{\mu}^{a})+\sqrt{\frac{\lambda}{\alpha}}\delta_{J(\varepsilon^{cd})}(\vartheta_{\mu}^{a}),$
@@
$:=$ $\delta_{J(\varepsilon^{cd})}(\frac{1}{2}\varpi_{\mu}^{ab})+\sqrt{\alpha\lambda}K^{ab}_{d,c}\vartheta_{\mu}^{d}\varepsilon^{c},$
where we denoted by $\varepsilon^{a}$ and $\varepsilon^{ab}$ the coordinate-dependent parameters of the gauge transformations [^5] Hereafter, $\delta$ without subscripts (GCT) or (CGCT) will always refer to gauge transformations, and their parameters, functions of spacetime coordinates, will be enclosed in parentheses. and we have singled out the different contributions along the generators of the Lie algebra. At the same time, the connection transforms infinitesimally under general coordinate transformation (GCT) $y^{\mu}(x)=x^{\mu}-\xi^{\mu}(x)$, according to
$\delta_{GCT(\xi)}\Omega^{I}_{\mu}=\mathcal{L}_{\xi}\Omega^{I}_{\mu}=\xi^{\nu}\partial_{\nu}\Omega_{\mu}^{I}+\Omega_{\nu}^{I}\partial_{\mu}\xi^{\nu}.$ (4.3)
@@
$\begin{split}\delta_{GCT(\xi)}\vartheta^{a}_{\mu}&=\xi^{\nu}\partial_{\nu}\vartheta_{\mu}^{a}+\vartheta_{\nu}^{a}\partial_{\mu}\xi^{\nu},\\ \delta_{GCT(\xi)}\varpi^{ab}_{\mu}&=\xi^{\nu}\partial_{\nu}\varpi_{\mu}^{ab}+\varpi_{\nu}^{ab}\partial_{\mu}\xi^{\nu}.\end{split}$ (4.4)
In order to study the interplay of gauge and coordinate transformations, it is convenient to consider a covariant generalization of the latter (CGCT) [Freedman_VanProeyen_2012]. For the full gauge potential this is defined as
@@
$\begin{split}\delta_{CGCT(\xi)}(\frac{1}{2}\varpi^{ab}_{\mu})&=\delta_{GCT(\xi)}(\frac{1}{2}\varpi^{ab}_{\mu})\\ &-\left(\partial_{\mu}(\xi^{\rho}\frac{1}{2}\varpi^{ab}_{\rho})+(\xi^{\rho}\frac{1}{2}\varpi^{cd}_{\rho})\frac{1}{2}\varpi^{ef}_{\mu}f^{ab}_{ef,cd}+(\xi^{\rho}\vartheta^{c}_{\rho})\vartheta^{d}_{\mu}\lambda K^{ab}_{d,c}\right),\end{split}$ (4.7)
@@
$\begin{split}\delta_{CGCT(\xi)}\vartheta^{a}_{\mu}&=\delta_{GCT(\xi)}\vartheta^{a}_{\mu}-\delta_{P(\xi^{\rho}\vartheta^{c}_{\rho})}(\vartheta_{\mu}^{a})-\delta_{J(\xi^{\rho}\varpi^{cd}_{\rho})}(\vartheta_{\mu}^{a}),\\ \delta_{CGCT(\xi)}\varpi^{ab}_{\mu}&=\delta_{GCT(\xi)}\varpi^{ab}_{\mu}-\delta_{J(\xi^{\rho}\frac{1}{2}\varpi^{cd}_{\rho})}(\varpi_{\mu}^{ab})-2\lambda\xi^{\rho}\vartheta^{c}_{\rho}\vartheta_{\mu}^{d}K_{d,c}^{ab}.\end{split}$ (4.8)
@@
$\begin{split}\delta_{CGCT(\xi)}\vartheta^{a}_{\mu}=\xi^{\rho}\mathfrak{T}^{a}_{\rho\mu},\end{split}$ (4.9)
@@
$\begin{split}\delta_{CGCT(\xi)}\varpi^{ab}_{\mu}=\xi^{\rho}\mathfrak{R}^{ab}_{\rho\mu}-2\lambda\xi^{\rho}\vartheta^{c}_{\rho}\vartheta_{\mu}^{d}K_{d,c}^{ab}.\end{split}$ (4.10)
@@
$\begin{split}\delta_{GCT(\xi)}\vartheta^{a}_{\mu}-\delta_{P(\xi^{\rho}\vartheta^{c}_{\rho})}(\vartheta_{\mu}^{a})-\delta_{J(\xi^{\rho}\varpi^{cd}_{\rho})}(\vartheta_{\mu}^{a})&=\xi^{\rho}\mathfrak{T}^{a}_{\rho\mu},\\ \delta_{GCT(\xi)}\varpi^{ab}_{\mu}-\delta_{J(\xi^{\rho}\frac{1}{2}\varpi^{cd}_{\rho})}(\varpi_{\mu}^{ab})&=\xi^{\rho}\mathfrak{R}^{ab}_{\rho\mu}.\end{split}$ (4.11)
As anticipated, CGCT allow one to rewrite the diffeomorphisms in terms of gauge transformations and covariant tensors. The first of Eqs. (4.11), appropriately rearranged, gives the soldering equation
@@
$\begin{cases}\vartheta^{a}\rightarrow\theta^{a}\,\,\,\text{tetrad one-form}\\ \varpi^{ab}\rightarrow\omega^{ab}\,\,\,\text{Lorentz connection one-form}\end{cases}$ (4.13)
Consequently, by recalling the definition of the fields in Eqs. (2.5) and (2.6), we can give a consistent geometric interpretation to the components of the curvature two-form $\mathcal{F}$:
[…]